# Kai Yao    Marc Juarez

### v1

@@ VI-D Utility–Robustness Trade-off
To examine the trade-off between attribution accuracy and adversarial robustness for MFD methods, we plot this relationship for both removal and forgery attacks in Figure 3 under white-box and black-box scenarios respectively. Broadly, we observe that the utility–robustness tension is pronounced for both removal and forgery attacks: methods with higher attribution accuracy tend to suffer higher ASR, indicating weaker robustness. Most high-utility methods (accuracy >0.7>0.7) exhibit significantly higher vulnerability than low-utility ones.
For example, neural network–based methods such as Wang20 achieve the highest attribution accuracy; yet, it is also among the most fragile methods to both removal and forgery, particularly under white-box attacks. Nonetheless, there are notable exceptions: for instance, Marra19a and Song24-RGB under black-box attacks, and Song24-Freq under white-box attacks. These methods pair moderate attribution accuracy with substantially stronger robustness, making them stand out as more balanced candidates among the evaluated approaches.
Overall, these findings highlight a fundamental trade-off: methods with high attribution accuracy often lack robustness, whereas more robust methods typically sacrifice accuracy. Promising directions for future work include: (i) improving residual- and manifold-based approaches such as Marra19a and Song24 to boost accuracy; (ii) strengthening CNN-based methods like Wang20 against adversarial perturbations; and (iii) exploring hybrid or ensemble designs that combine the robustness of residual/manifold methods with the strong attribution accuracy of CNN-based approaches.
@@ VI-E Attack Variance Across Models
Figure 4: Model vulnerability per MFD method, averaged over all applicable attack types. Each subplot plots removal ASR (x) vs. forgery ASR (y) for the generative models (indicated by colors and markers) tied to that method; ASR spans 0–1 and error bars show standard deviation across different attack types. The final subplot averages the data over all MFD methods (with standard deviation representing statistics over different MFD methods), showing general vulnerability trend of generative models - models with more resilience to removal are more vulnerable to forgery, and vice versa.
To examine whether particular generative models are inherently more vulnerable to certain adversarial attacks, we evaluated per-model vulnerability across all MFD methods. Figure 4 plots forgery ASR versus removal ASR for each model under each method, averaging over all applicable attack strategies. We define a model’s removal vulnerability as how easily its own images can be perturbed to be misclassified as coming from any other model, and its forgery vulnerability as how easily images from other models can be perturbed to be misclassified as coming from this model.
The results show that vulnerability is not consistently high or low for any given model or model family. Instead, it varies strongly with the MFD method. For example, in removal, NVAE is among the least vulnerable under McCloskey18 and Song24-Freq, yet appears highly vulnerable under other methods. Likewise, ADM is markedly more robust in Nowroozi22 than all other models, but not in any of the remaining methods. In forgery, ADM is significantly more vulnerable under Nataraj19 and Nowroozi22, yet this pattern does not generalize across all MFD methods.
Despite this variability, an important and obvious trend emerges: across MFD methods, image generation models that are more robust to removal tend to be more vulnerable to forgery, and vice versa. This trade-off is visible in the final subplot of Figure 4, and is also particularly clear under methods such as Nowroozi22, Song24-Freq, and Qian20. Intuitively, if a model’s fingerprint occupies a more isolated decision region, it may be easy for gradients to escape that region (removal) but harder to steer perturbations into it (forgery). We further observe that diffusion and VAE models are more resistant to removal and more vulnerable to forgery, whereas GAN-based models show the opposite trend. We leave a deeper investigation of these behaviors to future work.
@@ VII Discussion
Researchers have developed MFD techniques to provide provenance of model outputs and thus address some of the accountability challenges posed by today’s increasingly powerful generative models. In this paper, we provide the first systematic security analysis of these techniques, which includes over 14 MFD techniques across a range of possible threat models and attack strategies. Despite the reported high accuracy of these techniques in a clean setting in the literature, our findings reveal that they are highly vulnerable to adversarial manipulations.
Our results show that most techniques fail to provide reliable model attribution under removal attacks. For most MFD techniques, our attacks remove model fingerprints across various access assumptions without visually degrading the attacked images. Most concerning, simple black-box image perturbations succeed in removing fingerprints, with attacks achieving over $90\%$ success for some techniques. These findings indicate that fingerprint traces can be disrupted without sophisticated attacks or knowledge of the underlying MFD system.
Implications for MFD Technique Deployment. These results fundamentally undermine the applications proposed in the academic literature. The stark contrast between MFD’s high accuracy in the clean setting (often $90\%$) and its fragility under attack, uncovers a gap that can lead to a false sense of security if unacknowledged. For example, such high attack success rates imply that investigators and model providers cannot use these techniques to reliably trace images to the source models, compromising their forensic value and undermining the accountability they are intended to provide. Importantly, while forgery attacks are significantly less successful than removal, effective removal alone severely compromises MFD’s value for such applications.
Even in the clean setting, existing evaluations may have overestimated MFD performance for model attribution. Current studies evaluate small closed worlds of at most 20 different models, raising questions about how representative they are of real-world scenarios. Practical deployments could be far more complex. For example, detecting the source of an arbitrary image would require distinguishing among orders of magnitude more models. Furthermore, real-world scenarios might involve highly similar models: the ML community often reuses public datasets for training and relies on open sourced learning algorithms, potentially producing similar model fingerprints—a situation not represented by the small, diverse model pools considered in the literature. Conversely, other practical scenarios might involve smaller model pools. Additional sources of evidence can narrow down the candidate pool to one closer to those in the literature, potentially enabling effective MFD deployment. Therefore, we encourages future papers presenting new MFD schemes to evaluate them in realistic open worlds. However, the success of our attacks implies that these techniques remain unreliable even in such ideal conditions, highlighting the urgent need for developing MFD techniques that are robust in adversarial environments.
Implications for the Adversarial Arms-Race. Our results also have implications for the strategies chosen by attackers and fingerprinters. Since attackers cannot predict which MFD technique will be applied, they must develop attacks capable of removing fingerprints across all MFD techniques. This situation might seem like a whack-a-mole game that favors the fingerprinter, who can apply multiple techniques simultaneously. However, our results add nuances to this: we observe a tension between utility and robustness, where the most robust methods (e.g., Marra19a, Song24-RGB) exhibit higher detection error than the more vulnerable ones. Consequently, even when the fingerprinter deploys all 14 techniques, the attacks increase overall attribution uncertainty and thus undermine value of these techniques for forensic applications.
Limitations and Future Work. One of the main limitations of our work is that we cannot conclude that the removal attacks completely erase the fingerprints: they supress enough signal to impair model attribution under a specific experimental setting, but there may be enough signal to still enable deepfake detection. Future work could further investigate the effectiveness of the attacks against MFD techniques for deepfake detection.
Another avenue for future work is a more thorough exploration of the attack hyperparameter space. Although we tuned parameters that had significant impact, we made no attempt at optimization. Similarly, although future work could explore attacks tailored to specific techniques, attackers must be strategic: since attackers cannot predict which MFD techniques fingerprinters will deploy, attackers are better off increasing overall maximum success across techniques. Therefore, even highly successful bespoke attacks provide only marginal advantage if other techniques remain robust or if increased success against one method comes at the cost of reduced effectiveness against others.
From the perspective of mitigations, an important direction for future work is to evaluate defenses from adversarial machine learning, such as adversarial training or randomized smoothing, applied to the attribution model. A thorough investigation would require adapting these defenses to handle both removal and forgery attacks and developing corresponding adaptive attacks to properly assess robustness. Such methods must also consider the typical utility–robustness trade-off, as improved robustness often comes at the cost of reduced clean accuracy. We view this as a substantial research effort that merits a dedicated study, and thus leave the systematic design and evaluation of defenses for future work.
Additionally, we suspect that fingerprint suppression may introduce detectable patterns. We leave for future work investigating the detectability of the attacks. This capability would be particularly valuable for countering forgery attacks, as early detection could prevent adversaries from successfully misattributing content. Another avenue for defense is to explore hybrid approaches. In particular, despite their lower detection accuracies, our results identify Marra19a and Song24-RGB’s fingerprint features as particularly robust under black-box conditions. Future work could refine the features to maintain robustness while improving performance.
Active vs. Passive Provenance Mechanisms. Watermarking has emerged as an alternative to fingerprinting for providing similar accountability guarantees in the deployment of generative models. The increased control of watermarking over the identifying signal makes active techniques more suitable for adversarial environments. Recently proposed watermarking schemes demonstrate robustness against sophisticated removal attempts while providing undetectability and formal security guarantees [yang2024gaussian, gunn2024undetectable]. On the other hand, watermarking requires integration during model training or inference, which might introduce friction in its adoption by the industry. We encourage future work to further explore trade-offs between these two types of techniques or strengthen performance and robustness by combining the best of both worlds.
@@ VIII Conclusion
Model fingerprint detection (MFD) has been proposed as a means of providing provenance and accountability in the deployment of image generation models, with reported accuracies often exceeding $90\%$ in non-adversarial evaluations. However, our findings show that such results offer an incomplete picture of their reliability in the adversarial settings where attribution mechanisms are most critical. Through a systematic evaluation of 14 MFD techniques against 5 attack strategies, we expose a critical security gap: adversaries can reliably suppress or forge fingerprints, undermining the practical utility of current MFD methods for real-world attribution. While industry adoption of attribution mechanisms is already underway, it is unlikely to pause until robust solutions emerge. This urgency underscores the need not only to advance more resilient MFD techniques but also to develop complementary or alternative accountability mechanisms that remain effective under adversarial conditions. Looking forward, we identify several promising directions for strengthening model provenance of AI-generated images, including but not limited to leveraging more inherently robust methods such as residual- or manifold-based approaches, adversarially robust training of MFD systems, and hybrid designs that combine complementary strengths across existing model attribution techniques such as model watermarking.
@@ Acknowledgments
This work was supported by the Edinburgh International Data Facility (EIDF) and the Data-Driven Innovation Programme at the University of Edinburgh. Access to EIDF was facilitated through the University of Edinburgh’s Generative AI Laboratory GAIL Fellow scheme.
@@ Appendix A Image Quality Under Attacks
Example Images. As an example, Figure 5 provides a comprehensive visualization of successful removal attacks against the Song24-Freq MFD method across different generative models and attack strategies. The grid structure clearly illustrates how various attack types (W3, B1, and B2 with different perturbation techniques) perform on different target models, with each cell showing the original image, the perturbed result, and the scaled difference map. This visualization demonstrates the effectiveness of different attack strategies in removing model fingerprints while preserving the visual quality of the generated images. The difference images, scaled for visibility, reveal the spatial distribution of perturbations introduced by each attack method.
Figure 5: Visualization of successful removal attacks against the Song24-Freq MFD method for selected models. Each row represents a different generative model (target for fingerprint removal), and each column represents a different attack type. Within each cell, the three images show (from left to right): the original image, the perturbed image after attack, and the scaled difference. The difference is linearly mapped such that the lowest value corresponds to black and the highest value corresponds to white, with intermediate values scaled accordingly. The attack types include W3 (surrogate extractor), B1 (surrogate classifier), and B2 (basic image perturbations: blur, Gaussian noise, JPEG compression, and resize). The visualization demonstrates the effectiveness of various attack strategies in removing model fingerprints while maintaining visual quality. An “N/A” entry indicates non-existence of a successfully attacked image for the case.
@@ Appendix B Removal ASR vs Forgery ASR
Here in Figure 6 we report the relationship between removal-attack ASR and forgery-attack ASR for each gradient-based attack type separately, rather than aggregating them using the best white-box ASR as done in the main text (i.e. Figure 2). Specifically, we present results for W1, W2, and W3 individually. Note that certain attack types are only applicable to specific fingerprinting methods (e.g., W1 requires differentiability and is therefore only compatible with MFD methods such as Giudice21, Corvi23-R, Corvi23-S, Qian20, and Wang20). Importantly, the strong positive correlation between removal and forgery ASR persists for each attack type considered independently, reinforcing the robustness of the trend reported in the main paper.
@@ Appendix C Utility-Robustness Trade-off
Here in Figure 7 we report the trade-off relationship between utility and attack ASR for each gradient-based attack type separately, rather than aggregating them using the best white-box ASR as done in the main text (i.e. Figure 3). Specifically, we present results for W1, W2, and W3 individually. Note that certain attack types are only applicable to specific fingerprinting methods.
Figure 7: Trade-off between attribution utility (accuracy of the classifier $h$, y-axis) and robustness (ASR, x-axis). We plot the data with regard to each attack type separately in this figure, as opposed to Figure 3. Each point is an MFD method; error bars show standard deviation across 5 independent runs. Methods nearer the top-left deliver stronger utility–robustness trade-offs.
@@ Appendix D Hyperparameter Tuning
We conducted hyperparameter tuning for several attack cases with respect to their strength parameters to illustrate the trade-off between ASR and image quality degradation. Specifically, we report ablation results for multiple attack cases using MFD Dzanic20 as a representative example, as shown in Figures 8 and 9.
@@
Figures 8 and 9 plot ASR and image quality indicators (LPIPS/PSNR) as functions of these attack-strength hyperparameters. As expected, increasing attack strength generally yields higher ASR, but at the cost of noticeably degraded image quality, particularly for PGD-based attacks.

### v2

@@ VI-D Utility–Robustness Trade-off
To examine the trade-off between attribution accuracy and adversarial robustness for MFD methods, we plot in Figure 3 this relationship for both removal and forgery attacks under white- and black-box scenarios. Broadly, we observe a utility–robustness tension for both removal and forgery attacks: methods with higher attribution accuracy tend to suffer higher ASR, indicating weaker robustness. Most high-utility methods (accuracy >0.7>0.7) exhibit significantly higher vulnerability than low-utility ones.
For example, neural network–based methods such as Wang20 achieve the highest attribution accuracy but are also among the most vulnerable to both removal and forgery, particularly under white-box attacks. Nonetheless, there are notable exceptions: for instance, Marra19a and Song24-RGB under black-box attacks, and Song24-Freq under white-box attacks. These methods pair moderate attribution accuracy with substantially stronger robustness, making them stand out as the most balanced candidates among all the evaluated approaches.
Overall, these findings highlight a fundamental trade-off: methods with high attribution accuracy often lack robustness, whereas more robust methods typically sacrifice accuracy. The results indicate a few directions for future work, including: (i) improving residual- and manifold-based approaches such as Marra19a and Song24 to boost accuracy; (ii) strengthening CNN-based methods like Wang20 against adversarial perturbations; and (iii) exploring hybrid or ensemble designs that combine the robustness of residual/manifold methods with the strong attribution accuracy of CNN-based approaches.
@@ VI-E Attack Variance Across Models
Figure 4: Model vulnerability per MFD method, averaged over all applicable attack types. Each subplot shows removal ASR ($x$) vs. forgery ASR ($y$) for the generative models (indicated by colors and markers) tied to that method; error bars show standard deviation across different attack types. The final subplot averages the data over all MFD methods (with standard deviation representing statistics over different MFD methods), showing a general vulnerability trend of generative models: models with more resilience to removal are more vulnerable to forgery, and vice versa.
To examine whether particular generative models are inherently more vulnerable to certain adversarial attacks, we evaluated per-model vulnerability across all MFD methods. Figure 4 plots forgery ASR versus removal ASR for each model under each MFD method, averaging over all applicable attack strategies.
The results show that attack vulnerability is not consistently high or low for any given model or model family. Instead, it varies strongly with the MFD method. For example, in removal, NVAE is among the least vulnerable under McCloskey18 and Song24-Freq, yet appears highly vulnerable under other methods. Likewise, ADM is markedly more robust in Nowroozi22 than all other models, but not in any of the remaining methods. In forgery, ADM is significantly more vulnerable under Nataraj19 and Nowroozi22, yet this pattern does not generalize across all MFD methods.
Despite this variability, a trend emerges: across MFD methods, image generation models that are more robust to removal tend to be more vulnerable to forgery, and vice versa. This trade-off is visible in the final subplot of Figure 4, and is also particularly clear under methods such as Nowroozi22, Song24-Freq, and Qian20. Intuitively, if a model’s fingerprint occupies a more isolated decision region, it may be easy for gradients to escape that region (removal) but harder to steer perturbations into it (forgery). We further observe that diffusion and VAE models are more resistant to removal and more vulnerable to forgery, whereas GAN-based models show the opposite trend. We leave a deeper investigation of these behaviors to future work.
@@ VII Discussion
This paper asks a practical question: how well does model fingerprint detection (MFD) hold up when the content creator actively tries to evade attribution? We address this by evaluating 14 MFD methods on a 12-model attribution task under three attacker knowledge levels (White-box, Black-box Access I, and Black-box Access II), two attacker goals (removal and forgery), and five attack strategies.
Our results support three main conclusions. (1) Removal attacks are effective. For most MFD methods, the attacker can remove the fingerprints with imperceptible perturbations, even without information about the model or the MFD technique (Table II). (2) Forgery is relatively less effective but should not be dismissed. We still observe significant forgery success for many MFD methods, especially in white-box settings (Table III). (3) High clean accuracy does not imply robustness. Our results uncover a tension between clean attribution performance and adversarial robustness: none of the evaluated MFD methods is both consistently accurate and robust across the threat models we consider.
VII-A Implications for Provenance Use-cases
Many MFD papers motivate attribution as a provenance mechanism. Our findings suggest an important caveat: a method that performs well on clean benchmarks is unlikely to achieve the same performance in adversarial settings. This is especially relevant in forensic applications, where the subjects of an investigation have incentives to remove or forge fingerprints. Based on our evaluation, existing MFD methods are not yet ready for deployment in these contexts. A practical takeaway is that the evaluation methodology of MFD techniques in such applications should consider robustness, and should clearly define the threat model considered.
It is also important to evaluate attribution under more realistic candidate sets. Real-world deployments would face large, open-world scenarios and frequent model updates that introduce closely related variants, such as fine-tuned checkpoints. Studying how attacks scale to realistic candidate pools is essential for understanding the practical effectiveness of MFD techniques, both with and without adversarial interference.
VII-B What the Robustness Patterns Suggest
The patterns in Tables II–III indicate that robustness is largely method-driven. Some MFD approaches appear relatively more resistant in constrained black-box settings, especially residual- and manifold-based MFD methods. However, this robustness is not universal and may not hold under white-box access. A useful way to view this is that black-box robustness can reflect limited attack transferability under a particular perturbation budget, rather than a fundamental inability to manipulate the fingerprint. This distinction matters operationally: if the attribution pipeline is public, it has leaked, or can be closely approximated, the threat of adaptive attacks becomes more relevant.
These observations suggest a clear design objective for future research in MFD: methods should be developed and reported with an explicit robustness target, rather than focusing solely on clean accuracy. Based on the utility-robustness trade-off perspective, we identify two promising directions: (i) improving the clean accuracy of inherently robust representations, potentially leveraging NN-based classifiers to boost performance; and (ii) hardening high-accuracy fingerprints—such as MFD methods based on NN-learned features—against adaptive perturbations without sacrificing clean performance, potentially through techniques like adversarial training or randomized smoothing [cohen2019certified].
VII-C Limitations and Future Work
First, we examine multi-class attribution, a task inherently more complex than binary deepfake detection. A critical distinction lies in the relationship between attack types: while untargeted removal and targeted forgery are functionally equivalent in binary detection, they diverge significantly in the attribution setting. In multi-class attribution, untargeted removal is generally easier to achieve than evading detection of a deepfake because the adversary can exploit any of the multiple decision boundaries surrounding the correct class manifold. In contrast, targeted forgery remains more difficult, as it requires a precise shift into a specific, potentially distant region of the feature space. Consequently, an attack that successfully breaks attribution (removal) may still leave the sample within the broader manifold of synthetic content, failing to evade deepfake detection. Future work could study the gap between “detection-robustness” and “attribution-robustness” under the same threat model.
Second, while our attacks are strong and cover multiple access levels, they do not span the full space of adaptive strategies. Future work could explore broader perturbation families and multi-step post-processing pipelines, and should also study whether fingerprint suppression or forgery leaves detectable traces that could be used as a warning signal that a verifier can flag for a more careful inspection.
Third, the scope of our evaluation is constrained by the specific generator and dataset configuration (12 generators trained on FFHQ at $256\times 256$ resolution). Evaluating a broader range of generators and datasets, higher resolutions, and diverse types of distribution shifts would help quantify robustness under more diverse deployment conditions and thus provide a more complete understanding of MFD robustness.
Finally, our results suggest that provenance systems should not rely on a single mechanism alone. Watermarking can provide a strong and unambiguous provenance signal when a watermark is embedded at generation time and remains detectable, and in such cases it is natural to rely primarily on the watermark. However, watermarking is often unavailable in practice: many images are generated by open-source or legacy models without watermarking, detection may require secret keys or vendor-side access, and real-world distribution or editing pipelines can remove or break watermark detectability altogether. Passive fingerprinting addresses a different need. It does not require cooperation at generation time and is detected post hoc, which allows provenance analysis even when watermarking is unavailable or not possible. Fingerprinting is therefore not a universal replacement for watermarking, nor is watermarking a replacement for fingerprinting. It is thus worthwhile exploring hybrid provenance systems that use watermarking as a high-confidence signal when present, and falls back to fingerprinting when it is not, while explicitly accounting for removal, forgery, and uncertainty under adversarial manipulation [yang2024gaussian, gunn2024undetectable].
@@ VIII Conclusion
Model fingerprint detection has been proposed as a scalable means of providing provenance and accountability for AI-generated images, with many methods reporting strong attribution accuracy in non-adversarial evaluations. In this paper, we provide the first systematic security evaluation of such techniques for model attribution, formalizing three adversary knowledge levels and benchmarking 14 representative methods against five attack strategies under two attacker objectives: fingerprint removal and fingerprint forgery.
Our results reveal a critical gap between clean and adversarial performance. Fingerprint removal is highly effective for most methods through small, imperceptible perturbations, even in black-box regimes. Forgery, in contrast, is less effective but can still succeed in black-box settings and achieves high success under white-box access on some methods. We also identify a utility–robustness trade-off: techniques with the highest clean attribution accuracy are frequently among the most vulnerable, and no evaluated method achieves both high accuracy and high robustness across all threat models.
These findings suggest that current passive fingerprinting approaches alone are insufficient for high-stakes provenance in adversarial settings. Advancing reliable attribution will likely require (i) more robust fingerprint representations (e.g., residual/manifold-based), (ii) incorporating security considerations in the development and evaluation of new techniques, and (iii) hybrid designs that combine complementary provenance mechanisms, including active approaches like watermarking.
@@ Acknowledgments
We thank our anonymous reviewers for their valuable feedback, which substantially improved the paper. This work was supported by the Edinburgh International Data Facility (EIDF) and the Data-Driven Innovation Programme at the University of Edinburgh. Access to EIDF was facilitated through the University of Edinburgh’s Generative AI Laboratory GAIL Fellow scheme. Marc Juarez is a GAIL fellow and the recipient of a Google Research Scholar Award in Security.
@@ Appendix A Image Quality Under Attacks
Example Images. As an example, Figure 5 provides a comprehensive visualization of successful removal attacks against the Song24-Freq MFD method across different generative models and attack strategies. The grid structure clearly illustrates how various attack types (W3, B1, and B2 with different perturbation techniques) perform on different target models, with each cell showing the original image, the perturbed result, and the difference, which has been scaled to improve visibility. This visualization demonstrates the effectiveness of different attack strategies in removing model fingerprints while preserving the visual quality of the generated images. The difference images show the spatial distribution of perturbations introduced by each attack.
Figure 5: Visualization of successful removal attacks against the Song24-Freq MFD method for selected models. Each row represents a different generative model (target for fingerprint removal), and each column represents a different attack type. Within each cell, the three images show (from left to right): the original image, the perturbed image after attack, and the scaled difference. The difference is linearly mapped such that the lowest value corresponds to black and the highest value corresponds to white, with intermediate values scaled accordingly. The attack types include W3 (surrogate extractor), B1 (surrogate classifier), and B2 (basic image perturbations: blur, Gaussian noise, JPEG compression, and resize). The visualization demonstrates the effectiveness of various attack strategies in removing model fingerprints while maintaining visual quality. An “N/A” value indicates that there is no successfully attacked image for that entry.
@@ Appendix B Removal ASR vs Forgery ASR
In Figure 6 we report the relationship between removal ASR and forgery ASR for each gradient-based attack type separately, rather than aggregating them using the best white-box ASR as done in the main text (Figure 2). Specifically, we present results for W1, W2, and W3 individually. Note that certain attack types are only applicable to specific fingerprinting methods (e.g., W1 requires differentiability and is therefore only compatible with MFD methods such as Giudice21, Corvi23-R, Corvi23-S, Qian20, and Wang20). Importantly, the strong positive correlation between removal and forgery ASR persists for each independent attack type, reinforcing the trend reported in the main text.
@@ Appendix C Utility-Robustness Trade-off
In Figure 7 we report the trade-off between utility and attack ASR for each gradient-based attack type separately, rather than aggregating them using the best white-box ASR as done in the main text (Figure 3). Specifically, we present results for W1, W2, and W3 individually. Note that certain attack types are only applicable to specific fingerprinting methods.
Figure 7: Trade-off between attribution utility (accuracy of classifier $h$, $y$-axis) and robustness (ASR, $x$-axis). We plot the data with regard to each attack type separately, as opposed to Figure 3. Each point is an MFD method; error bars show standard deviation across 5 independent runs. Methods nearer the top-left achieve a better compromise between utility and robustness.
@@ Appendix D Hyperparameter Tuning
We conducted hyperparameter tuning for several attack cases with respect to their strength parameters to illustrate the trade-off between ASR and image quality degradation. Specifically, we report ablation results for multiple attack cases using Dzanic20 as a representative example, as shown in Figures 8 and 9.
@@
Figures 8 and 9 plot ASR and image quality indicators (LPIPS/PSNR) as functions of these hyperparameters. As expected, increasing attack strength generally yields higher ASR, but at the cost of noticeably degraded image quality, particularly for PGD-based attacks.
Figure 8: Effect of perturbation strength (PGD’s $\epsilon$) on attack success and image quality for removal (W3, B1) and forgery (B1) attacks. The curves show that there is a trade-off between robustness and image quality regulated by perturbation strength.